# Reasoning Runtime Behavior of a Program with LLM: How Far Are We?

### v2

@@ III-C1 Description of Incremental Consistency
According to the above descriptions, we find that the knowledge required to finish the previous task is contained by that of the following task. Intuitively, the subsequent tasks are more difficult than the current one. Hence, if a model fails to correctly complete a task (e.g., fails to finish CPP) but then predicts the following tasks correctly (e.g., correct prediction of output), we consider this model behaves inconsistently in consecutive tasks.
@@ III-C2 Evaluation Approach
In addition, depending on how many consecutive times consistency is maintained (i.e., $\{1,1,1,1\}$ means 4 times), we assign different weights to reward models that maintain Incremental Consistency more often. It is intuitive because it is harder to behave incrementally consistently across more sub-tasks.
Finally, we define Incremental Consistency Score (IC Score) to quantitatively model IC of an LLM $\mathcal{M}$. For our benchmark that contains $N$ number of $(\mathbf{P},\mathbf{X},I,V)$ pairs, ICS can be computed as:
$\text{ICS}=\frac{100}{N}\sum_{i=1}^{N}\text{ICS}_{i},$
$\text{ICS}_{i}=\begin{cases}\dfrac{1}{2^{j-1}},&\text{if }\mathbf{R}_{i}=\mathbf{S}_{j},j\in\{1,2,3,4\}\\ 0,&\text{otherwise}\end{cases}$
@@
  2. 2.
    If the answers are partially correct, and Incremental Consistency is observed, the model gains a partial score.
  3. 3.
@@ III-D Benchmark Construction
Our framework utilizes existing executable benchmarks to evaluate the code reasoning ability of LLMs, and we introduce how to adapt these base benchmarks into our framework in two steps: (i) Runtime Behavior Extraction; (ii) Problem Construction and Filtering:
@@ III-D2 Problem Construction and Filtering
CCP and EPP. In these two tasks, we focus on whether and when a statement is being executed. As the execution sequence of statements could be very long for the structures such as the loop, we analyze the control flow graph and break the program into several blocks. We prioritize the last statement in the block which is generally more difficult than statements that are executed sequentially.
PSP. The program state might be complicated and contains various types of statements and variables (e.g., built-in variables of Python). In this paper, we inspect the code and mainly pay attention to the following types of statements: (i) Assignment. We extract the variable(s) at the left-hand side for assignment statements. The possible types of variables are identifier (i.e., ordinary variables like “x”), subscript (i.e., array slices like “x[0]”), and attribute (i.e., fields like “x.y”). In most cases, we extract identifiers. Note that some naive assignments such as a = 0 or l = [] are skipped, but we keep statements like “a += 1” for the change of variable value; (ii) Return. In return statements, we extract the variables in the returned object if local variables are returned. If the returned object is a constant value, we will select the “nearest” variable, i.e., the last variable that is not constant. (iii) Others. For other situations, if some variables after the current line are changed, we extract the changed variables and obey the priority of “new variable $>$ changed variables $>$ changed attributes (i.e., self.xxx)”. We ignore some objects that are hard to express and compare such as “self” object.
OP. In output prediction task, we follow CRUXEval [14] and utilize the assertion statements in the test cases of the base benchmarks. For base benchmarks like ClassEval [22], where one test case contains multiple assertions, we use all the assertions. Specifically, we replace the right operands in assertions with question marks (“??”), and challenge models to predict the masked values.
@@ IV-A Base Benchmarks
In our experiments, we first need to obtain the runtime behavior of code such as program state, thus base benchmarks should be executable and equipped with test cases. Moreover, we would like to experiment with diverse types of data (e.g., different programming scenarios). We utilize existing code benchmarks as the basis for code reasoning evaluation. Specifically, we select the following two popular benchmarks:
@@
These benchmarks are widely used in previous research [15, 42, 43]. We show the statistics of the adapted benchmarks in Table I.
@@
Number of Problems 3152
# Avg. Tokens in Program 14.0
# Avg. Tokens in Selected Statement 408.3
IV-B Studied Code LLMs
As a result, considering these dimensions we select several state-of-the-art code LLMs that have been applied to various code related tasks [46, 47, 48]. Table II presents detailed features of them and we can see that full consideration of the diversity of models across various features is taken to enhance the generalization of our study.
@@ IV-C Prompt Design
In our work, we utilize prompting to evaluate code LLMs with our code reasoning tasks. We refer a recent study [14] on code reasoning and design our prompt templates as illustrated in Figure 3. For classic few-shot prompting, our prompt template mainly consists of five parts including system message, few-shot demonstrations, program, question, and answer. If Chain-of-Thought (CoT) [49] prompting is utilized, the thoughts are added into the prompt template as well as the examples in it. Both few-shot prompting and CoT prompting are widely applied in various tasks [50, 49], including reasoning tasks [14].
@@ IV-D Implementation Details
Access of Models and Base Benchmarks. For open-source models such as CodeLlama, we use the corresponding official releases available on HuggingFace [51]. For closed-source models (e.g., GPT-3.5-Turbo and GPT-4-Turbo), we invoke the OpenAI API [52] to access them. Two benchmarks (i.e., HumanEval and ClassEval) are also publicly available on HuggingFace [53, 54].
@@
Configurations. Temperature can control the randomness in the generated results of models [8]. Specifically, we follow Gu et al. [14] and set the temperature to 0.8. For tasks with direct prompting, we set the maximum length of generated tokens to 256, while for tasks with CoT prompting, we set it to 1024. For the rest of the parameters, we use the default settings in vLLM, to ensure a fair comparison. To obtain reliable results, experiments for all open-source models with few-shot prompting are repeated five times, and we report the mean and standard deviations of numbers in Section V. We do not repeat experiments for closed-source models like GPT series due to a limited budget.
V Results
  • RQ1: How do LLMs perform on Runtime Behavior Reasoning?
  • RQ2: How do LLMs perform on Incremental Consistency evaluation?
V-A RQ1: Performance of Runtime Behavior Reasoning
@@
Table III shows the detailed results of Runtime Behavior Reasoning. All the models are evaluated with few-shot prompting except for special annotation (i.e., CodeLlama-7b-Instruct (CoT)). Below, we discuss the results from different aspects.
@@
Overall Performance. Overall, we find that the performance of different LLMs presents a large variation, and GPT-4-Turbo shows superior performance in reasoning about program execution. For example, GPT-4-Turbo achieves the best results in all metrics of Runtime Behavior Reasoning, and its average accuracy outperforms the second best (i.e., 55.7% of GPT-3.5) by a large margin (i.e., an absolute improvement of 19.3%). However, the overall performance of open-source models is not high, and the best performer among them (i.e., CodeLlama-34B-Instruct) only achieves a level close to that of GPT-3.5 (i.e., 51.0% v.s. 55.7%) in terms of average accuracy.
@@
Task. Runtime Behavior Reasoning consists of four distinct evaluation tasks, i.e., CCP, PSP, EPP, and OP, and the performance varies among different tasks. For example, all models achieve an accuracy of over 50% in OP (i.e., Output Prediction), while only about half of them (i.e., 8 out of 15) can provide the correct answers for more than 10% problems in EPP (i.e., Execution Path Prediction). Hence, according to the average score of tasks, the performance distribution may suggest that EPP is the most challenging task and OP is relatively easy among them.
@@
Size and Category. In general, we observe that for models within the same family, the variant with larger size of parameters shows better performance in Runtime Behavior Reasoning. In the case of CodeLlama-instruct series, as the number of parameters increases (i.e., from 7B to 34B), the accuracy of EPP has a relative improve by over 100% (i.e., 14.4% → 29.2%). Meanwhile, smaller models such as StarCoder2-3B can also outperform larger models like CodeLlama-7B-Instruct in terms of average accuracy (i.e., 37.3% v.s. 35.8%). The StarCoder2 series utilizes varied architectures and training datasets compared to CodeLlama. This may demonstrate that apart from the size of parameters, the model architecture and training strategy also play an important role in the ability of code reasoning. We also find that code LLMs do not exhibit an obviously leading advantage over general LLMs with the same size.
@@
Training Strategy. As shown in Table III, we conduct experiments on three variants of CodeLlama (i.e., base, instruct, and Python) with the same 7B size. Compared to the base model, the instruct variant that leverages instruction tuning techniques brings gains in the code reasoning ability (i.e., Avg. Acc. from 35.8% to 38.5%), which may reflect the relationship between understanding instructions and reasoning program. Meanwhile, since our base benchmarks are all in Python, additional training with Python corpora (i.e., CodeLlama-Python) leads to an improvement of performance (i.e., a 0.9% absolute improvement). In addition, we notice that although further fine-tuning applied to Magicoder series improves their performance on code generation [7], the improvement of code reasoning ability is relatively limited compared to their foundation model CodeLlama-7B-Python. This may indicate that the training strategies they utilize are not well-suited for reasoning tasks in our evaluation.
@@
Prompting Strategy. The last row of Table III presents the performance of CodeLlama-7B-Instruct with CoT prompting. Compared to the model with few-shot prompting, the performance of CPP, PSP, and EPP receives varying degrees of improvement. For instance, the EPP accuracy with CoT is improved from 10.8% to 21.4%, surpassing the performance of the larger 13B model (i.e., 14.4%). It may demonstrate the effectiveness of presenting how to reason about a piece of code step by step. Nevertheless, CoT prompting fails to improve its OP performance, with an absolute decrease of 6.8%. This may result from the wrong thought chain that the model generates for the whole program, and eventually leads to the mistake.
@@
Summary for RQ1: Models with different features (e.g., size and training strategy) exhibit notable disparities in performance on Runtime Behavior Reasoning. Overall, GPT-4-Turbo demonstrates a clear advantage over other models across all four tasks in our setting.
@@ V-B RQ2: Incremental Consistency Evaluation
Figure 4: The combination figure includes the bar chart for the average accuracy of Runtime Behavior Reasoning and the line graph for the score of Incremental Consistency.
Figure 4 shows the sorted average accuracy for Runtime Behavior Reason for different models, with an additional line indicating the IC scores. A detailed information of mean IC scores and their standard deviations are also reported in Table III.
Overall Performance. We find that the majority of LLMs exhibit a low level of Incremental Consistency with scores below 20, which highlights the inconsistency in model behavior across the four tasks of Runtime Behavior Reasoning. Among all of the models, GPT-4-Turbo stands out with the highest IC score of 42.5, even more than double that of the second place GPT-3.5 (i.e., 20.6). Given that GPT-4-Turbo achieves the best results in both Runtime Behavior Reasoning and Incremental Consistency Evaluation (i.e., 75.0% of Avg. Acc. and 42.5 of IC score), we believe that it has both superior ability on program reasoning and a high level of Incremental Consistency across sequentially related tasks in our evaluation.
Trend Between IC and Runtime Behavior Reasoning. As illustrated in Figure 4, we find that there is an approximately similar trend between the model’s average accuracy and its IC score. For example, Compared to CodeLlama-7B-Instruct, its larger version CodeLlama-34B-Instruct has a noticeable higher average accuracy (i.e., 38.5% v.s. 51.0%) and IC score (i.e., 4.1 v.s. 11.8). However, this pattern does not hold for all models. The performance of the general LLM Mistral-7B is not as good as that of CodeLlama-34B in terms of average accuracy (i.e., 48.0% v.s. 51.0% in Runtime Beheavior Reasoning), but it performs better on Incremental Consistency (i.e., 16.3 v.s. 11.8).
Others. Similar to the results of RQ1, we observe that code LLMs do not significantly outperform general LLMs in IC evaluation. For example, the IC score of Gemma-2B-It (i.e., 5.5) is higher than code LLMs trained with more code corpora like CodeLlama-7b-Instruct and StarCoder2-7B (i.e., 4.1 and 4.2). This phenomenon may suggest that more code data cannot help LLMs reason programs better and maintain their Incremental Consistency. In addition, CoT prompting for CodeLlama-7B-Instruct leads to a great increase in its IC score (i.e., from 4.1 to 7.5), and this improvement of Incremental Consistency may benefit from the explicit problem-solving steps.
Summary for RQ2: In code reasoning tasks, most LLMs behave inconsistently and their average accuracy is not entirely associated with IC. GPT-4-Turbo achieves an IC score as high as 42.5, surpassing other models by a large margin (i.e., more than 21.9 absolute improvements).

### v3

@@ III-C1 Description of Incremental Consistency
According to the above descriptions, we find that the knowledge required to finish the previous task is contained by that of the following task. Intuitively, the subsequent tasks are more difficult than the current task. Hence, if a model fails to correctly complete a task (e.g., fails to finish CPP) but then predicts the following tasks correctly (e.g., correct prediction of output), we consider this model behaves inconsistently in consecutive tasks.
@@ III-C2 Evaluation Approach
[…]
In addition, depending on how many consecutive times consistency is maintained (i.e. $\{1,1,1,1\}$ means 4 times), we assign different weights to reward models that maintain Incremental Consistency more often. It is intuitive because it is harder to behave incrementally consistently across more sub-tasks.
Finally, we define Incremental Consistency Score (IC Score) to quantitatively model the Incremental Consistency of an LLM $\mathcal{M}$. For our benchmark that contains $N$ number of $(\mathbf{P},\mathbf{X},I,V)$ pairs, IC Score can be computed as:
IC $\text{IC Score}=\frac{100}{N}\sum_{i=1}^{N}\text{IC Score}_{i},$
$\text{IC Score}_{i}=\begin{cases}\dfrac{1}{2^{j-1}},&\text{if }\mathbf{R}_{i}=\mathbf{S}_{j},j\in\{1,2,3,4\}\\ 0,&\text{otherwise}\end{cases}$
@@
  2. ii)
    If the answers are partially correct and Incremental Consistency is observed, the model gains a partial score.
  3. iii)
@@ III-D Benchmark Construction
As Fig. 2 illustrates, our framework utilizes existing executable benchmarks to evaluate the code reasoning ability of LLMs. We introduce how to adapt these base benchmarks into our framework in two steps: (i) Runtime Behavior Extraction; (ii) Problem Construction and Filtering:
@@ III-D2 Problem Construction and Filtering
CCP and EPP. In these two tasks, we focus on whether and when a statement is being executed. As the actual execution sequence of statements could be very long for loops, we analyze the control flow graph and break the program into several blocks. We prioritize the last statement in a block as it leads to various new blocks and is generally more difficult to reason about.
PSP. This task challenges an LLM to predict the type and value of a variable. For $\bm{\mathcal{R}}$Eval, we inspect the code and focus on the following types of statements:
  1. i)
    Assignment. We extract the variable(s) at the left-hand side for assignment statements. The possible types of variables are identifier (i.e., ordinary variables like “x”), subscript (i.e., array slices like “x[0]”), and attribute (i.e., fields like “x.y”). In most cases, we extract identifiers. Note that some naive assignments such as a = 0 or l = [] are skipped, but we keep statements like a += 1 for the change of variable value.
  2. ii)
    Return statement. In return statements, we extract the variables in the returned object if local variables are returned. If the returned object is a constant value, we will select the “nearest” variable, i.e., the last variable that is not constant.
  3. iii)
    Others. For other situations, if any variable after the current line is changed, we extract a changed variable based on the priority of “new variable $>$ changed variables $>$ changed attributes (i.e., self.xxx)”. Not all variables are used because we prefer variables that have closer logical relationship with other tasks (e.g., EPP). We ignore objects of non-serializable classes or complex structures (e.g., “self” objects), as it is challenging to convert them to canonical string representations and compare ground truth with an LLM’s output.
OP. In output prediction task, we follow CRUXEval [14] and utilize the assertion statements in the test cases of the base benchmarks. For base benchmarks such as ClassEval [22], where one test case contains multiple assertions, we use all the assertions. Specifically, we replace the right operands in the assertions with question marks (“??”), and challenge models to predict the masked values.
@@ IV-A Base Benchmarks
In our experiments, we first need to obtain the runtime behavior of code such as program state, thus base benchmarks should be executable and equipped with test cases. Moreover, we would like to experiment with diverse types of data (e.g., different programming scenarios). Therefore, we utilize existing code generation benchmarks as the basis for code reasoning evaluation. Typically, code generation benchmarks can be categorized into two types: competition-level ones (i.e., with standalone functions) [46, 9] and context-aware ones (i.e., with more code context like class and file dependencies) [22, 47]. Considering the diversity of base benchmarks, we choose HumanEval as a representative of competition-level benchmarks and ClassEval as a representative of context-aware benchmarks.
@@
These benchmarks are widely used in previous research [15, 48, 49]. We show the statistics of the adapted benchmarks in Table I. In addition to our selection, we highlight that our framework is applicable for other similar code generation benchmarks.
@@
# of Problems 3152
# of Avg. Tokens in Programs 408.3
# of Avg. Tokens in Selected Statements 14.0
@@
IV-B Studied Code LLMs
As a result, considering these dimensions we select several state-of-the-art code LLMs that have been applied to various code related tasks [52, 53, 54, 29]. Table II presents detailed features of them and we can see that full consideration of the diversity of models across various features is taken to enhance the generalization of our study.
@@ IV-C Prompt Design
In our work, we utilize prompting to evaluate code LLMs with our code reasoning tasks. We refer to a recent study [14] on code reasoning and design our prompt templates as illustrated in Fig. 3. For classic few-shot prompting, our prompt template consists mainly of five parts, including the system message, few-shot demonstrations, program, question, and answer. If Chain-of-Thought (CoT) [55] prompting is utilized, the thoughts are added into the prompt template as well as the examples in it. Both few-shot prompting and CoT prompting are widely applied in various tasks [56, 57, 55], including reasoning tasks [14].
@@ IV-D Implementation Details
Access of Models and Base Benchmarks. For open-source models such as CodeLlama, we use the corresponding official releases available on HuggingFace [58]. For closed-source models (e.g., GPT-3.5-Turbo and GPT-4-Turbo), we invoke the OpenAI API [59] to access them. Two benchmarks (i.e., HumanEval and ClassEval) are also publicly available on HuggingFace [60, 61]. To help replicate our research, we list detailed information, such as model IDs and URLs in our replication package [62].
@@
Configurations. Temperature can control the randomness in the generated results of models [8]. Specifically, we follow Gu et al. [14] and set the temperature to 0.8. For tasks with direct prompting, we set the maximum length of generated tokens to 256, while for tasks with CoT prompting, we set it to 1024. For the rest of the parameters, we use the default settings in vLLM, to ensure a fair comparison. To obtain reliable results, experiments for all open-source models with few-shot prompting are repeated five times, and we report the mean and standard deviation values in Section V. We do not repeat experiments for closed-source models like the GPT series due to a limited budget.
@@
V Results
  • RQ1: How do LLMs perform on Runtime Behavior Reasoning?
  • RQ2: How do LLMs perform on Incremental Consistency evaluation?
V-A RQ1: Performance of Runtime Behavior Reasoning
@@
Table III shows the detailed results of Runtime Behavior Reasoning. All models are evaluated with few-shot prompting except for special annotated ones (i.e., CodeLlama-7b-Instruct (CoT)). Below, we discuss the results from different aspects.
@@
Overall Performance. Overall, we find that the performance of different LLMs presents a large variation, and GPT-4-Turbo shows superior performance in reasoning about program execution. For example, GPT-4-Turbo achieves the best results in all metrics of Runtime Behavior Reasoning, and its average accuracy outperforms the second best (i.e., 55.7% of GPT-3.5) by a large margin (i.e., an absolute improvement of 19.3%). However, the overall performance of open-source models is not high and the best performer among them (i.e., CodeLlama-34B-Instruct) only achieves a level close to that of GPT-3.5 (i.e., 51.0% v.s. 55.7%) in terms of average accuracy.
@@
Task. Runtime Behavior Reasoning consists of four distinct evaluation tasks, i.e., CCP, PSP, EPP, and OP, and the performance varies among different tasks. For example, all models achieve an accuracy of more than 50% in OP (i.e., Output Prediction), while only about half of them (i.e., 8 out of 15) can provide the correct answers for more than 10% problems in EPP (i.e., Execution Path Prediction). Hence, according to the average score of tasks, the performance distribution may suggest that EPP is the most challenging task and OP is relatively easy among them.
@@
Size and Category. In general, we observe that for models within the same family, the variant with larger size of parameters shows better performance in Runtime Behavior Reasoning. In the case of the CodeLlama-instruct series, as the number of parameters increases (i.e., 7B $\rightarrow$ 34B), the accuracy of EPP has a relative improve by over 100% (i.e., 14.4% $\rightarrow$ 29.2%). Meanwhile, smaller models like StarCoder2-3B can also outperform larger models such as CodeLlama-7B-Instruct in terms of average accuracy (i.e., 37.3% v.s. 35.8%). The StarCoder2 series utilizes varied architectures and training datasets compared to CodeLlama. This may demonstrate that apart from parameter size, the model architecture and training strategy also play an important role in code reasoning ability. We also find that code LLMs do not exhibit an obviously leading advantage over general LLMs of the same size.
@@
Training Strategy. As shown in Table III, we conduct experiments on three variants of CodeLlama (i.e., base, instruct, and Python) of the same 7B size. Compared to the base model, the “instruct” variant that leverages instruction tuning techniques brings gains in the code reasoning ability (i.e., Avg. Acc. from 35.8% to 38.5%), which may reflect the relationship between understanding instructions and reasoning program. Meanwhile, since our base benchmarks are all in Python, additional training with Python corpora (i.e., CodeLlama-Python) leads to an improvement of performance (i.e., an absolute improvement of 0.9%). In addition, we note that although further fine-tuning applied to the Magicoder series improves their performance in code generation [7], the improvement in code reasoning ability is relatively limited compared to their foundation model CodeLlama-7B-Python. This may indicate that the training strategies they utilize are not well suited for the reasoning tasks in our evaluation.
@@
Prompting Strategy. The last row of Table III presents the performance of CodeLlama-7B-Instruct with CoT prompting. Compared to the model with few-shot prompting, the performance of CPP, PSP, and EPP receives varying degrees of improvement. For instance, the EPP accuracy with CoT is improved from 10.8% to 21.4%, surpassing the performance of the larger 13B model (i.e., 14.4%). It may demonstrate the effectiveness of presenting how to reason about a piece of code step by step. However, CoT prompting fails to improve its OP performance, with an absolute decrease of 6.8%. This may result from the wrong thought chain that the model generates for the whole program and eventually leads to the mistake.
@@
Summary for RQ1: Models with different features (e.g., size and training strategy) exhibit notable disparities in performance on Runtime Behavior Reasoning. Overall, GPT-4-Turbo demonstrates a clear advantage over other models in all four tasks in our setting.
@@ V-B RQ2: Incremental Consistency Evaluation
Figure 4: Average Accuracy of Runtime Behavior Reasoning and Incremental Consistency Score for different models, sorted in descending order according to average accuracy.
Fig. 4 shows the sorted average accuracy for Runtime Behavior Reason for different models, with an additional line indicating the IC scores. A detailed information of mean IC scores and their standard deviations are also reported in Table III.
Overall Performance. We find that the majority of LLMs exhibit a low level of Incremental Consistency with scores below 20, which highlights the inconsistency in model behavior across the four tasks of Runtime Behavior Reasoning. Among all of the models, GPT-4-Turbo stands out with the highest IC score of 42.5, even more than double that of the second place GPT-3.5 (i.e., 20.6). Given that GPT-4-Turbo achieves the best results in both Runtime Behavior Reasoning and Incremental Consistency Evaluation (i.e., 75.0% of Avg. Acc. and 42.5 of IC score), we believe that it has both superior ability in program reasoning and a high level of Incremental Consistency across sequentially related tasks in our evaluation.
Trend Between IC and Runtime Behavior Reasoning. As illustrated in Fig. 4, we find that there is an approximately similar trend between the model’s average accuracy and its IC score. For example, compared to CodeLlama-7B-Instruct, its larger version CodeLlama-34B-Instruct has a noticeable higher average accuracy (i.e., 38.5% v.s. 51.0%) and IC score (i.e., 4.1 v.s. 11.8). However, this pattern does not hold for all models. The performance of the general LLM Mistral-7B is not as good as that of CodeLlama-34B in terms of average accuracy (i.e., 48.0% v.s. 51.0% in Runtime Beheavior Reasoning), but performs better on Incremental Consistency (i.e., 16.3 v.s. 11.8).
Others. Similar to the results of RQ1, we observe that code LLMs do not significantly outperform general LLMs in the IC evaluation. For example, the IC score of Gemma-2B-It (i.e., 5.5) is higher than code LLMs trained with more code corpora like CodeLlama-7b-Instruct and StarCoder2-7B (i.e., 4.1 and 4.2). This phenomenon may suggest that more code data cannot help LLMs reason programs better and maintain their Incremental Consistency. In addition, CoT prompting for CodeLlama-7B-Instruct leads to a great increase in its IC Score (i.e., 4.1 $\rightarrow$ 7.5), and this improvement of IC Score may benefit from explicit problem solving steps.
Summary for RQ2: In code reasoning tasks, most LLMs behave inconsistently and their average accuracy is not entirely associated with IC. GPT-4-Turbo achieves an IC Score of as high as 42.5, surpassing other models by a large margin (i.e., more than 21.9 absolute improvements).